# Confidence Aware Learning for Reliable Face Anti-spoofing

### v1

@@ II-C Cross-domain FAS methods
Another commonly used approach to improve the FAS generalization is through domain generalization (DG) methods. The DG-based methods utilize multiple datasets simultaneously to enhance the generalization of FAS models. Most DG-based FAS works adopt the idea of feature alignment [16, 18, 20], making the FAS feature indistinguishable across multiple training datasets. Subsequent works introduce the disentangle methods, selecting task-irrelevant features, such as identity features [23, 26] or style features [25], and disentangle them from the FAS feature to reduce the bias. Additionally, meta-learning is also an important research direction in FAS domain generalization. Different training datasets are used as meta-train sets and meta-test sets separately, thereby learning effective model structures [27] and FAS patterns [30]. A recent study encourages the model to converge towards an optimal flat minimum [36], achieving great results. However, due to the limitations of attack types in the training data, these DG-based methods still can never perform well on all the unknown presentation attacks.
@@ IV-C Effectiveness of the MD-based Confidence
The results indicate that the proposed output-based prediction confidence also helps the models understand their capability boundaries to some extent. Both “Baseline” and SSDG achieve performance improvements on HKBU-MARs and ROSE-Youtu by filtering out uncertain samples. However, we also observe that these two models do not obtain consistent improvement with increasing thresholds on SiW and WMCA, showing even worse performance on samples with higher confidence. This suggests that the output-based confidence is not stable in various test conditions.
The proposed CA-FAS using Mahalanobis distance-based confidence generally performs more stable and better across different test datasets. We find that CA-FAS performs relatively poorly on HKBU-MARs. We believe this is because HKBU-MARs only includes spoof data with the unseen 3D mask attacks. Therefore, even after filtering out a certain proportion of low confidence samples using the quantile threshold, the remaining samples with 3D mask attacks still lie beyond the model’s capability boundary. In contrast, for other test datasets that contain at least one type of known presentation attacks, the proposed model’s performance significantly improves after applying the quantile threshold. The results align perfectly with our expectations. Next, we will further analyze the reasons behind these differences.
@@ IV-E Ablation Study
We conduct ablation experiments to test the performance of CA-FAS under different feature distribution constraints, and the results are shown in Table III. First, we try using only one Gaussian distribution, referred to as “one-class”. We find that under this setting, the model’s performance on various test datasets is suboptimal. We believe this is because clustering all features together reduces the distinction between live and spoof features, making it difficult for the classifier to differentiate between live and spoof samples.
Next, we test a setting similar to SSDG, where the training spoof samples are divided into different Gaussian distributions based on the domain, referred to as “domain-based”. We find that under this setting, the model performs well on SiW, indicating that SSDG’s domain-based division method helps with the cross-domain generalization of the features. However, we also observe that the model’s performance declines on the other three datasets containing unknown attacks. This suggests that while the domain-based approach aids in handling domain variations, it is less effective when dealing with unknown attack types.
Finally, we divide all spoof data into a single Gaussian distribution, resulting in a binary setup with two Gaussian distributions: one for live and one for spoof. This setting, labeled as “binary”, performs well on datasets with unknown mask attacks such as HKBU-MARs and WMCA. However, on the remaining two datasets, while the overall performance on the entire dataset ($p=0.0$) is acceptable, the model’s ability to assess its own prediction confidence is poor, leading to a decrease in performance after filtering out low confidence samples. At the bottom of the table, “attack-based” refers to the implementation of the feature constraints we use in the proposed CA-FAS, which demonstrate the most stable performance in the experiments.

### v2

@@ II-C Cross-domain FAS methods
Another commonly used approach to improve the FAS generalization is through domain generalization (DG) methods. The DG-based methods utilize multiple datasets simultaneously to enhance the generalization of FAS models. Most DG-based FAS works adopt the idea of feature alignment [16, 18, 20], making the FAS feature indistinguishable across multiple training datasets. Subsequent works introduce the disentangle methods, selecting task-irrelevant features, such as identity features [23, 26] or style features [25], and disentangle them from the FAS feature to reduce the bias. Additionally, meta-learning is also an important research direction in FAS domain generalization. Different training datasets are used as meta-train sets and meta-test sets separately, thereby learning effective model structures [27] and FAS patterns [30]. A recent study encourages the model to converge towards an optimal flat minimum [36], achieving great results. Another recent work further explores the use of the hierarchical relations in samples to enhance model’s generalization ability across unseen domains [37]. Manipulation of the test-time data style is also considered a promising approach for boosting FAS performance [38]. However, due to the limitations of attack types in the training data, these DG-based methods still can never perform well on all the unknown presentation attacks.
@@ IV-C Effectiveness of the MD-based Confidence
The results indicate that the proposed output-based prediction confidence also helps the models understand their capability boundaries to some extent. Both “Baseline” and SSDG achieve performance improvements on HKBU-MARs and ROSE-Youtu by filtering out uncertain samples. The “Baseline” with the output-based confidence even outperforms the proposed CA-FAS on HKBU-MARs across different values of $p$. This is because the model’s output tends to be label-overfitted under the cross-entropy optimization, making the output-based confidence more effective in evaluating the uncertainty introduced by the unknown 3D mask attacks in HKBU-MARs. However, we also observe that “Baseline” and SSDG do not show consistent improvement with increasing thresholds on SiW and WMCA, where unknown domains play a more significant role in making the model’s predictions unreliable. This suggests that the output-based confidence is not stable in all the test conditions.
In contrast, the Mahalanobis distance-based confidence in CA-FAS provides greater stability in unknown scenarios and attacks, consistently enhancing the model’s prediction reliability across all four test datasets by filtering out low-confidence samples.
@@ IV-E Ablation Study
We conduct ablation experiments to evaluate the performance of CA-FAS using different approaches to modeling the reliable range, and the results are shown in Table III. Specifically, we divide the training Gaussians from different perspectives, such as categorizing by different domains or distinguishing between live vs. spoof samples, and examine the FAS performance under these conditions.
First, we try using a single Gaussian distribution to fit all the training data, referred to as “one-class”. We find that under this setting, the model’s performance on various test datasets is suboptimal. We believe this is because clustering all features together reduces the distinction between live and spoof features, making it difficult for the classifier to differentiate between live and spoof samples.
Next, we test a setting similar to SSDG, where the training spoof samples are divided into different Gaussians based on their domain labels, while the training live samples are modeled using a single Gaussian, referred to as “domain-based”. We find that under this setting, the model performs well on SiW, indicating that SSDG’s domain-based division method helps with the cross-domain generalization of the features. However, we also observe that the model’s performance declines on the other three datasets containing unknown attacks. This suggests that while the domain-based approach aids in handling domain variations, it is less effective when dealing with unknown attack types.
Finally, we cluster the training live samples into a single Gaussian, while the training spoof samples are grouped into another Gaussian. This setting, labeled as “binary”, performs well on datasets with unknown mask attacks such as HKBU-MARs and WMCA. However, on the remaining two datasets, while the overall performance on the entire dataset ($p=0.0$) is acceptable, the model’s ability to assess its own prediction confidence is poor, leading to a decrease in performance after filtering out low confidence samples.
At the bottom of the table, “attack-based” refers to the implementation of the feature constraints employed in our proposed CA-FAS, which demonstrate the most stable performance in the experiments. We believe this is due to the fact that whether an input’s pattern resembles any known attack is the key factor in determining the reliability of the model’s prediction. Therefore, modeling the reliable range based on known attacks is the most effective method for assessing the model’s confidence.